3/5 = [ ]/20 3/5 = [ ]/20 3/5 = [ ]/20 3/5 = [ ]/20

 3/5 = [ ]/20
 3/5 = [ ]/20
 3/5 = [ ]/20
 3/5 = [ ]/20

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the missing number be y.
Then, (3/5) = (y/20)
By cross multiplication we get,
(3 × 20)/5 = y

y = 60/5
Divide both numerator and denominator by 5. y = 12
Therefore, 3/5 = [12]/20

Thanks